OpenStudy (anonymous):

y=5x+1 4x+y=10 help me Im desperate

jimthompson5910 (jim_thompson5910):

4x+y=10 4x+5x+1=10 ... plug in y = 5x + 1 (ie replace y with 5x+1) 9x + 1 = 10 9x = 10-1 I'll let you finish

jimthompson5910 (jim_thompson5910):

you first solve that for x then you use the solution for x to get y

OpenStudy (anonymous):

Can you help me? I have to do a bunch of these for extra credit in order to compete in sports next weekend and I do not know how to do this at all

jimthompson5910 (jim_thompson5910):

when you solve 9x = 10-1 for x, what do you get

OpenStudy (anonymous):

Uhhhhhhh I know you deffinitly get a number? Ha.....ha....ha o.o im so screwewd

jimthompson5910 (jim_thompson5910):

9x = 10-1 9x = 9 x = ???

OpenStudy (anonymous):

oh oh oh oh 1!!!

jimthompson5910 (jim_thompson5910):

don't panic and just take things one step at a time you'll be fine

jimthompson5910 (jim_thompson5910):

yep x = 1

OpenStudy (anonymous):

:D

jimthompson5910 (jim_thompson5910):

now we know that y = 5x + 1 so... y = 5x + 1 y = 5(1) + 1 y = ???

OpenStudy (anonymous):

uhhhhh 6?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so together, the solution is x = 1, y = 6

jimthompson5910 (jim_thompson5910):

they form the ordered pair (1,6)

OpenStudy (anonymous):

So does that make it (1,6)?

OpenStudy (anonymous):

Oh yay! :D I like love you right now haha

jimthompson5910 (jim_thompson5910):

what this means is that if you graphed both equations they would be two lines that cross at the point (1,6)

jimthompson5910 (jim_thompson5910):

yes you got it

OpenStudy (anonymous):

yayayayayayaya haha now I finnially get it o.o I see the light!

jimthompson5910 (jim_thompson5910):

lol that's good, you need it to read and such

OpenStudy (anonymous):

Lol yea light is important